A dragster, starting from rest, accelerates at 28 m/s/s for 4.5 s. Then the parachute comes out and it slows to 46 m/s in 10 s. What is its acceleration in those 10 s when the parachute is deployed
1
Expert's answer
2020-10-23T11:50:06-0400

Let's first find the velocity of the dragster just before the parachute comes out:


"v=v_0+at,"

here, "v" is the velocity of the dragster just before the parachute comes out, "v_0=0\ \dfrac{m}{s}" is the initial velocity of the dragster, "a=28\ \dfrac{m}{s^2}" is the acceleration of the dragster during time "t=4.5\ s."

Then, from this equation we can find "v":


"v=at=28\ \dfrac{m}{s^2}\cdot 4.5\ s=126\ \dfrac{m}{s}."

Finally, we can find the acceleration of the draster in those 10 s when the parachute is deployed from the same kinematic equation:


"v=v_0+at,"

here, "v=46\ \dfrac{m}{s}" is the final velocity of the dragster when the parachute is deployed, "v_0=126\ \dfrac{m}{s}" is the initial velocity of the dragster just before the parachute comes out, "t=10\ s" is the time during which the dragster slows down.

Then, from this equation we can find the acceleration of the draster in those 10 s when the parachute is deployed:


"a=\dfrac{v-v_0}{t}=\dfrac{46\ \dfrac{m}{s}-126\ \dfrac{m}{s}}{10\ s}=-8\ \dfrac{m}{s^2}."

The sign minus indicates that the dragster decelerates.

Answer:

"a=-8\ \dfrac{m}{s^2}."